A 45 grams of copper is heated from 35 ᵒC to 85 ᵒC.

How much heat is absorbed by the copper ? The specific heat of copper is 0.093.



1
Expert's answer
2021-05-21T17:06:14-0400

"0.093\ (kcal/(kg\cdot K))=389.4\ (J/(kg\cdot K))"


"Q=cm\Delta T=389.4\cdot0.045\cdot(85-35)=876.15\ (J)"